OpenStudy (anonymous):

simplify ((sqrt)98n^18)

OpenStudy (anonymous):

\[\sqrt{98n ^{18}}\]

OpenStudy (mathstudent55):

Can you factor 98 into its prime factors?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Can you write n^18 as the square of something?

OpenStudy (anonymous):

so 98 an 1 2 and 49

OpenStudy (anonymous):

7 and 14

OpenStudy (mathstudent55):

Correct, but you need something square. 7 and 14 doesn't help. 2 and 49 helps. 2 * 7 * 7 = 2 * 49 49 can be written as the square of something. What is it?

OpenStudy (anonymous):

2401 is 46^2 but the square root of 49=7

OpenStudy (mathstudent55):

46^2 is not 4601, but what does that have to do with this problem?

OpenStudy (anonymous):

noting soory i dindt know waht u wanted from me

OpenStudy (mathstudent55):

Ok, you have 98 = 2 * 7^2 Now can you write n^18 as something square?

OpenStudy (mathstudent55):

Remember the rule (a^b)^c = a^(bc) For example, (5^3)^6 = 5^(3*6) = 5^18

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

9!

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

still here? @mathstudent55

OpenStudy (mathstudent55):

great n^18 = (n^9)^2

OpenStudy (mathstudent55):